The reactant concentration in a first-order reaction was 8.00×10−2 M after 15.0 s and 5.40×10−3 M after 85.0 s . What is the rate constant for this reaction?
Answer #1

K1 = 1/tln(a/a-x)

(1/15)ln(a/0.08) = (1/85)ln(a/0.0054)

a = initial concentration of reactant = 0.1425 M


k1 = (1/15)ln(0.1425/0.08) = 0.0385 s-1